OpenStudy (anonymous):

factor x^2-4x-12 in the form of x^2+bx+c

OpenStudy (anonymous):

This is expanded form, not factored: x^2+bx+c o_O

mathslover (mathslover):

hello @PatriciaMae first of all welcome to openstudy : \[\huge{x^2-4x-12}\] \[\huge{x^2+bx+c}\] b = -4 and c = 12 it is already inthe form of x^2+bx+c

OpenStudy (anonymous):

It can however be factored: Let: 1 = a -6 + 2 = -4 = b -6 * 2 = -12 = c So therefore: \[(x-6) (x+2)\] Check it by putting back in expanded form: (x-6) (x+2) F . O . I . L . (x)(x) + (x)(2) + (-6)(x) + (-6)(2)

OpenStudy (unklerhaukus):

to factor an expression in the form \[ax^2+bx+c\] use the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

@UnkleRhaukus , that works for all values of x for quadratic equations, but I do feel: 1. Factoring is easier and faster here. :-) (personal opinion) 2. What she wrote was in expression not an equation, the quadratic equation needs: ax^2 + bx + c = 0 If you need to solve for when that whole quadratic polynomial is equal to zero as an equation, then either factoring or the quadratic equation works here: (x-6)(x+2)=0 x-6=0 -or- x+2=0 x=6 -or- x=-2